State the hypotheses and identify the claim, find the critical value(s), compute the test value, make the decision, summarize the results (make the appropriate statement of the results of the claim included rejection or non-rejection of the null hypothesis). Please show all of your work.

Four clubs were randomly selected, and quantity of their members in the clubs are noted. At = 0.01, can it be concluded that there is a dependent relationship between the clubs and the political party affiliation of its members?

Club 1 Club 2 Club 3 Club 4

Democrat 33 22 18 13

Republican 13 16 10 23
Answer #1

hypothesis:-

H_{0}: the relationship between the clubs and the political party affiliation of its members are independent.

H1 the relationship between the clubs and the political party affiliation of its members are dependent [ claim ]

the critical value = 11.345

[ df = (4-1)*(2-1) = 3

chi square critical value for alpha= 0.01, df = 3, right tailed test = 11.345 ]

the necessary calculation table :-

club 1 club 2 club 3 club 4 row total
democrat observed 33 22 18 13 86
expected (46*86)/148 =26.7297 22.0811 16.2703 20.9189
\frac{(obs-exp)^2}{exp} 1.4709 0.0003 0.1839 2.9977
republican observed 13 16 10 23 62
expected 19.2703 15.9189 11.7297 15.0811
\frac{(obs-exp)^2}{exp} 2.0403 0.0004 0.2551 4.1581
column total 46 38 28 36 148

the test statistic be:-

\chi^2=\sum \frac{(obs-exp)^2}{exp}\approx \mathbf{11.107}

decision:-

\chi^2=11.107 < 11.345

we fail to reject the null hypothesis.

conclusion:-

There is not enough evidence to conclude that there is a dependent relationship between the clubs and the political party affiliation of its members.

*** if you have any doubts about the problem please mention it in the comment box... if satisfied please LIKE.